The accompanying table shows results of challenged referee calls in a major tennis tournament. Use a 0.05 significance level to test the claim that the gender of the tennis player is independent of whether a call is overturned. Was the Challenge to the Call​ Successful?

yes no
Men 297 848
Women 457 649

Determine the null and alternative hypotheses

Determine the​ P-value of the test statistic.

Use a 0.05 significance level to test the claim that the gender of the tennis player is independent of whether a call is overturned. Choose the correct answer below.
Answer #1

Null hypothesis : Ho : gender of the tennis player is independent of whether a call is overturned

Alternate Hypothesis : H1 : gender of the tennis player is dependent of whether a call is overturned

(0 Test Statistic Xstat

O : Observed Frequency

E: Expected Frequency

Row Total x Column Total Grand Total

Given,

O: Observed Frequency:

Was the Challenge to the Call​ Successful
yes no
Men 297 848
Women 457 649

Expected frequency: E

Was the Challenge to the Call​ Successful
yes no Total
Men (1145*754)/2251 (1145*1497)/2251 1145
Women (1106*754)/2251 (1106*1497)/2251 1106
Total 754 1497 2251
Was the Challenge to the Call​ Successful
yes no Total
Men 383.5318 761.4682 1145
Women 370.4682 735.5318 1106
Total 754 1497 2251
O E O-E (O-E)2 (O-E)2/E
297 383.5318 -86.5318 7487.7461 19.5231
457 370.4682 86.53176 7487.7461 20.2116
848 761.4682 86.53176 7487.7461 9.8333
649 735.5318 -86.5318 7487.7461 10.1800
Total 59.7481

Test Statistic Xstat

p - value P( > xat)= P(x > 59.7481 stat.

Degrees of freedom = (Number of rows - 1) * (Number of columns - 1) = (2-1)*(2-1)=1

For '1' degrees of freedom ; P(x259.7481) 0.000

P-value of the test statistic = 0.000

significance level : \small \alpha = 0.05 ;

As p-value: 0.000 < \small \alpha : 0.05; Reject the null hypothesis;

There is sufficient evidence to conclude that the gender of the tennis player is dependent of whether a call is overturned